### v2

# A coupled-channel perspective analysis on bottom-strange molecular pentaquarks

Abstract
In the present work, we systematically study various bottom-strange molecular pentaquarks to search for possible bound states and resonances by adopting a one boson exchange model within the complex scaling method. According to our calculations, we predict several bound and resonant states for bottom baryon $Y_{b}(\Lambda_{b},\Sigma_{b})\bar{K}^{(*)}$ and $Y_{b}K^{(*)}$ systems. In particular, a bound state in the $I(J^{P})=1/2(1/2^{-})\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ system may correspond to the particle $\Xi_{b}(6227)$. Meanwhile, the predicted bound state with mass in the range of $6303\sim 6269~\rm{MeV}$ in the $I(J^{P})=1/2(1/2^{-})\Sigma_{b}K/\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ system is flavor exotic and does not appear in the spectroscopy of conventional baryons, which may help clarify the nature of $\Xi_{b}(6227)$. We hope that our study can provide useful guidance for future experimental searches.
molecular states, coupled-channel analysis, complex scaling method
pacs:
12.39.Pn, 13.75.Lb, 14.40.Rt
I introduction
After the discovery of the $X(3872)$ by the Belle experiment Belle:2003nnu in 2003, a large amount of data has been accumulated over the past two decades in high energy collision experiments. Meanwhile, a series of new phenomenological studies related to the $XYZ$ and $P_{c}/T_{cc}$ states have been reported Liu:2019zoy ; Liu:2015fea . A detailed investigation of those exotic hadron states provides new insights into decoding their internal structures, which may deepen our understanding of the nonperturbative properties of quantum chromodynamics (QCD). Since many new particles are located near the hadron-hadron thresholds, these states could be naturally interpreted as candidates for molecular states Chen:2022asf ; Dong:2017gaw ; Guo:2017jvc ; Karliner:2017qhf ; Zou:2021sha ; Mai:2022eur ; Zou:2013af . It is highly desirable to identify those molecular states out of the numerous candidates and predict more of them for experimental searches, which will motivate the experimental search for such molecular states.
In 2021, the LHCb Collaboration reported two resonances, namely $X_{0}(2900)$ and $X_{1}(2900)$, in the $D^{-}K^{+}$ invariant mass spectrum by analyzing the decay amplitude of the $B^{+}\to D^{+}D^{-}K^{+}$ decay channel LHCb:2020pxc ; LHCb:2020bls . Since these two states are located near the $\bar{D}^{*}K^{*}$ and $\bar{D}_{1}^{*}K^{*}$ thresholds, they are regarded as hadronic molecule candidates Chen:2020aos ; He:2020btl ; Burns:2020epm ; Agaev:2020nrc ; Xiao:2020ltm ; Kong:2021ohg ; Ke:2022ocs . Recently, in the analysis of the $D_{s}^{+}\pi^{+}$ and $D_{s}^{+}\pi^{-}$ invariant mass spectra, the LHCb Collaboration has observed two new peaks, $T_{c\bar{s}0}^{0}(2900)$ and $T_{c\bar{s}0}^{++}(2900)$, whose masses and widths are $M(T_{c\bar{s}0}^{0})=2892\pm 14\pm 15$ $\rm MeV$, $\Gamma=119\pm 26\pm 12$ $\rm MeV$ and $M(T_{c\bar{s}0}^{++})=2921\pm 17\pm 19$ $\rm MeV$, $\Gamma=137\pm 32\pm 14$ $\rm MeV$ LHCb:2022sfr ; LHCb:2022lzp , respectively. Given their near-threshold behaviors and quantum numbers, these two $T_{c\bar{s}0}^{0(++)}$ states are proposed as isovector $D^{*}K^{*}$ molecules with $J^{P}=0^{+}$ Agaev:2022eyk ; Wang:2023hpp ; An:2022vtg .
Until now, most molecular candidates were observed in the charm sector, while the experimental observations in the bottom sector are still scarce. In 2006, the DØ Collaboration announced a narrow structure, referred to as the $X(5568)$, in the $B_{s}^{0}\pi^{\pm}$ channel D0:2016mwd . Then, the LHCb Collaboration investigated the $B_{s}^{0}\pi^{\pm}$ invariant mass spectrum, but no significant signal was found LHCb:2016dxl . Later, the ATLAS, CDF, and CMS Collaborations ATLAS:2018udc ; CDF:2017dwr ; CMS:2017hfy released similar results. Meanwhile, the $X(5568)$ has been theoretically discussed in previous works Chen:2016ypj ; Xiao:2016mho ; Agaev:2016urs ; Guo:2016nhb ; Albaladejo:2016eps ; Albaladejo:2016lbb , and cannot be assigned as an isovector $BK$ molecular state. In 2021, the LHCb Collaboration reported two states in the $B^{\pm}K^{\pm}$ mass spectrum, which are referred to as $B_{sJ}(6063)$ and $B_{sJ}(6114)$. If the missing photon from the $B^{*\pm}\to B^{\pm}\gamma$ was taken into consideration, the masses and widths were measured to be $B_{sJ}(6109):M=6108.8\pm 1.1\pm 0.7~\rm MeV$ and $\Gamma=22\pm 5\pm 4~\rm MeV$, and $B_{sJ}(6158):M=6158\pm 4\pm 5~\rm MeV$ and $\Gamma=72\pm 18\pm 25~\rm MeV$ LHCb:2020pet , respectively. In theory, the $B_{sJ}(6158)$ has been widely investigated in the literature Kong:2021ohg . Some of the existing works suggested that $B_{sJ}(6158)$ can be interpreted as a $\bar{B}K^{*}$ molecular state with $I(J^{PC})=0(1^{+})$. Also, several works showed that the existence of $\bar{B}^{(*)}K^{(*)}(B^{(*)}\bar{K}^{(*)})$ molecular states is allowed Kolomeitsev:2003ac ; Guo:2006fu ; Guo:2006rp ; Sun:2018zqs .
It can be seen that numerous exotic hadronic molecular states containing heavy quarks have been observed experimentally. The heavy quark symmetry is supposed to have been proven to play a significant role in predicting undiscovered states and understanding their production mechanisms, which has stimulated several theoretical studies Asanuma:2023atv ; Tanaka:2024siw ; Wang:2023eng ; Sakai:2023syt . In a previous work, the authors investigated the open-charm molecular counterparts of the newly observed $T_{c\bar{s}}^{a0(++)}$ composed of $Y_{c}(\Lambda_{c},\Sigma_{c})$ baryons and strange mesons $K^{(*)}$, by adopting the one-boson exchange model. From their estimations, some bound states may exist that correspond to the newly observed $T_{c\bar{s}}^{a0(++)}$ Chen:2022svh . According to heavy quark symmetry, in the bottom sector, the light diquark in heavy baryons $\Sigma_{b}/\Lambda_{b}$ has the same color structure as a light antiquark $\bar{q}$, as shown in Figure 1. If the $B_{sJ}(6158)$ can be explained as a $\bar{B}K^{*}$ molecular state with $I(J^{PC})=0(1^{+})$, one should also expect the existence of isoscalar $\bar{B}^{(*)}K^{(*)}$ molecular state. Under these circumstances, it is natural to conjecture whether open-bottom molecular pentaquarks could also exist. Moreover, it is worth mentioning that, in 2018, the LHCb Collaboration reported a peak in both the $\Lambda_{b}^{0}K^{-}$ and $\Xi_{b}^{0}\pi^{-}$ invariant mass spectra, named $\Xi_{b}(6227)$ LHCb:2018vuc . However, until now, whether the $\Xi_{b}(6227)$ should be accommodated into traditional baryon $\lambda$-mode $P$-wave $\Xi_{b}^{\prime}$ baryon with $J^{P}=3/2^{-}$ or $5/2^{-}$ Chen:2018orb ; He:2021xrh ; Wang:2018fjm ; Cui:2019dzj , or instead as a pure molecular state $\Sigma_{b}\bar{K}$ with $J^{P}=1/2^{-}$ Huang:2018bed ; Zhu:2020lza ; Mutuk:2024elj ; Ozdem:2021vry is still on discussion. Thus, it is both urgent and necessary to explore the possibility of $\Xi_{b}(6227)$ being a molecular state and to predict more bottom-strange molecular pentaquark candidates for future experimental searches.
Figure 1: A sketch of heavy superflavor symmetry between $\Sigma_{b}(\Lambda_{b})K^{(*)}$ pentaquarks and $\bar{B}K^{(*)}$ tetraquarks. $q$ stands for the light quarks ($u$ or $d$).
Recently, we systematically studied the hidden bottom molecular tetraquark with complex scaling method by adopting one-boson-exchange (OBE) model Song:2024ngu . In the present work, utilizing the same formalism, we systematically study various bottom-strange molecular pentaquarks to search for possible bound states and resonances by adopting the complex scaling method Aguilar:1971ve ; Balslev:1971vb ; Moiseyev:1998gjp ; Ho:1983lwa and Gaussian expansion method Hiyama:2003cu ; Hiyama:2018ivm . For bottom baryon $Y_{b}(\Lambda_{b},\Sigma_{b})$ and anti-strange meson $\bar{K}^{(*)}$ interactions, our calculations demonstrate that some bound and resonant states are revealed. For instance, in the $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ system, we obtain a bound state below $\Sigma_{b}\bar{K}$ threshold that can be regarded as the particle $\Xi_{b}(6227)$. Meanwhile, we extend our study to $Y_{b}K^{(*)}$ systems, and find two flavor exotic bound states, which can be searched for in future experiments.
The rest of this paper is organized as follows. We briefly introduce the formalism of effective interactions and complex scaling method in Sec. II. In Sec. III, we present the numerical results and discussions for the $Y_{b}$$K^{(*)}$ and $Y_{b}$$\bar{K}^{(*)}$ systems. Finally, we summarize in Sec. 5.
II Formalism of effective interaction and complex scaling method
II.1 The effective interactions
In this work, we adopt the one-boson-exchange model to describe the interactions between the hadrons and analyze the formation mechanisms of molecular states. The chiral symmetric interacting Lagrangian, which corresponds to the coupling between a bottom baryon and a light meson, can be constructed as Liu:2011xc
$\mathcal{L}_{\mathcal{B}_{\bar{3}}}$ $=$ $l_{B}\langle\bar{\mathcal{B}}_{\bar{3}}\sigma\mathcal{B}_{\bar{3}}\rangle+i\beta_{B}\langle\bar{\mathcal{B}}_{\bar{3}}v^{\mu}(\mathcal{V}_{\mu}-\rho_{\mu})\mathcal{B}_{\bar{3}}\rangle,$ (1)
$\mathcal{L}_{\mathcal{B}_{6}}$ $=$ $l_{S}\langle\bar{\mathcal{S}}_{\mu}\sigma\mathcal{S}^{\mu}\rangle-\frac{3}{2}g_{1}\varepsilon^{\mu\nu\lambda\kappa}v_{\kappa}\langle\bar{\mathcal{S}}_{\mu}A_{\nu}\mathcal{S}_{\lambda}\rangle$ (2)
$+$ $i\beta_{S}\langle\bar{\mathcal{S}}_{\mu}v_{\alpha}\left(\mathcal{V}_{ab}^{\alpha}-\rho_{ab}^{\alpha}\right)\mathcal{S}^{\mu}\rangle+\lambda_{S}\langle\bar{\mathcal{S}}_{\mu}F^{\mu\nu}(\rho)\mathcal{S}_{\nu}\rangle,$~
$\mathcal{L}_{\mathcal{B}_{\bar{3}}\mathcal{B}_{6}}$ $=$ $ig_{4}\langle\bar{\mathcal{S}^{\mu}}A_{\mu}\mathcal{B}_{\bar{3}}\rangle+i\lambda_{I}\varepsilon^{\mu\nu\lambda\kappa}v_{\mu}\langle\bar{\mathcal{S}}_{\nu}F_{\lambda\kappa}\mathcal{B}_{\bar{3}}\rangle+h.c..$ (3)
The axial current $A_{\mu}$, the vector current $\mathcal{V}_{\mu}$, and the vector meson field strength tensor $F^{\mu\nu}(\rho)$ are defined as follows
$\mathcal{V}_{\mu}$ $=$ $\frac{1}{2}(\xi^{{\dagger}}\partial_{\mu}\xi+\xi\partial_{\mu}\xi^{{\dagger}}),$ (4)
$A_{\mu}$ $=$ $\frac{1}{2}(\xi^{{\dagger}}\partial_{\mu}\xi-\xi\partial_{\mu}\xi^{{\dagger}}),$ (5)
$F_{\mu\nu}(\rho)$ $=$ $\partial_{\mu}\rho_{\nu}-\partial_{\nu}\rho_{\mu}+[\rho_{\mu},\rho_{\nu}],$ (6)
respectively. Here, $\xi=\text{exp}({P}/f_{\pi})$ and $\rho_{ba}^{\mu}=ig_{V}{V}_{ba}^{\mu}/\sqrt{2}$. The $\mathcal{B}_{\bar{3}}$, $\mathcal{B}_{\mu}=-\sqrt{\frac{1}{3}}(\gamma_{\mu}+\nu_{\mu})\gamma^{5}\mathcal{B}_{6}+\mathcal{B}_{6\mu}^{*}$, P{P}, and V{V} denote the matrices of the ground states of singly heavy baryon multiplets in $\bar{3}_{F}$, $6_{F}$, light pseudoscalar, and vector mesons, respectively, whose explicit forms read
$\mathcal{B}_{6}$ $=$ $\begin{pmatrix}\Sigma_{b}^{+}&\frac{\Sigma_{b}^{0}}{\sqrt{2}}&\frac{\Xi_{b}^{(^{\prime},*)0}}{\sqrt{2}}\\ \frac{\Sigma_{b}^{0}}{\sqrt{2}}&\Sigma_{b}^{-}&\frac{\Xi_{b}^{(^{\prime},*)-}}{\sqrt{2}}\\ \frac{\Xi_{b}^{(^{\prime},*)0}}{\sqrt{2}}&\frac{\Xi_{b}^{(^{\prime},*)-}}{\sqrt{2}}&\Omega_{b}^{(*)-}\end{pmatrix},\quad\mathcal{B}_{\bar{3}}=\begin{pmatrix}0&\Lambda_{b}^{0}&\Xi_{b}^{0}\\ -\Lambda_{b}^{0}&0&\Xi_{b}^{-}\\ -\Xi_{b}^{0}&-\Xi_{b}^{-}&0\\ \end{pmatrix}$, (7)
$P$ $=$ $\begin{pmatrix}\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta}{\sqrt{6}}&\pi^{+}&K^{+}\\ \pi^{-}&-\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta}{\sqrt{6}}&K^{0}\\ K^{-}&\bar{K}^{0}&\frac{2}{\sqrt{6}}\eta\\ \end{pmatrix}$$\quad$ (8)
$V=\begin{pmatrix}\frac{\rho^{0}}{\sqrt{2}}+\frac{\omega}{\sqrt{2}}&\rho^{+}&K^{*+}\\ \rho^{-}&-\frac{\rho^{0}}{\sqrt{2}}+\frac{\omega}{\sqrt{2}}&K^{*0}\\ K^{*-}&\bar{K}^{*0}&\phi\\ \end{pmatrix}$. (9)
Under the SU(3) symmetry, the effective Lagrangians describing the interactions between the strange mesons and light mesons can be expressed as Lin:1999ad
$\mathcal{L}_{PPV}$ $=$ $\frac{ig}{2\sqrt{2}}\langle\partial^{\mu}P\left(PV_{\mu}-V_{\mu}P\right\rangle,$ (10)
$\mathcal{L}_{VVP}$ $=$ $\frac{g_{VVP}}{\sqrt{2}}\epsilon^{\mu\nu\alpha\beta}\left\langle\partial_{\mu}V_{\nu}\partial_{\alpha}V_{\beta}P\right\rangle,$ (11)
$\mathcal{L}_{VVV}$ $=$ $\frac{ig}{2\sqrt{2}}\langle\partial^{\mu}V^{\nu}\left(V_{\mu}V_{\nu}-V_{\nu}V_{\mu}\right)\rangle.$ (12)
More specifically, one can further write the effective Lagrangian depicting the couplings as
$\mathcal{L}_{\sigma}$ $=$ $l_{B}\langle\bar{\mathcal{B}}_{\bar{3}}\sigma\mathcal{B}_{\bar{3}}\rangle-l_{S}\langle\bar{\mathcal{B}}_{6}\sigma\mathcal{B}_{6}\rangle,$ (13)
$\mathcal{L}_{{P}}$ $=$ $i\frac{g_{1}}{2f_{\pi}}\varepsilon^{\mu\nu\lambda\kappa}v_{\kappa}\langle\bar{\mathcal{B}}_{6}\gamma_{\mu}\gamma_{\lambda}\partial_{\nu}{P}\mathcal{B}_{6}\rangle$ (14)
- $\sqrt{\frac{1}{3}}\frac{g_{4}}{f_{\pi}}\langle\bar{\mathcal{B}}_{6}\gamma^{5}\left(\gamma^{\mu}+v^{\mu}\right)\partial_{\mu}{P}\mathcal{B}_{\bar{3}}\rangle+h.c.,$
$\mathcal{L}_{{V}}$ $=$ $\frac{1}{\sqrt{2}}\beta_{B}g_{V}\langle\bar{\mathcal{B}}_{\bar{3}}v\cdot{V}\mathcal{B}_{\bar{3}}\rangle-\frac{\beta_{S}g_{V}}{\sqrt{2}}\langle\bar{\mathcal{B}}_{6}v\cdot{V}\mathcal{B}_{6}\rangle$ (15)
- $\frac{\lambda_{I}g_{V}}{\sqrt{6}}\varepsilon^{\mu\nu\lambda\kappa}v_{\mu}\langle\bar{\mathcal{B}}_{6}\gamma^{5}\gamma_{\nu}\left(\partial_{\lambda}{V}_{\kappa}-\partial_{\kappa}{V}_{\lambda}\right)\mathcal{B}_{\bar{3}}\rangle+h.c.$
- $i\frac{\lambda g_{V}}{3\sqrt{2}}\langle\bar{\mathcal{B}}_{6}\gamma_{\mu}\gamma_{\nu}\left(\partial^{\mu}{V}^{\nu}-\partial^{\nu}{V}^{\mu}\right)\mathcal{B}_{6}\rangle,$
$\mathcal{L}_{K^{(*)}K^{(*)}\sigma}$ $=$ $g_{\sigma}m_{K}\bar{K}K\sigma-g_{\sigma}m_{K^{*}}\bar{K}^{*}\cdot K^{*}\sigma,$ (16)
$\mathcal{L}_{PKK^{*}}$ $=$ $\frac{ig}{4}\left[\left(\bar{K}^{*\mu}K-\bar{K}K^{*\mu}\right)\left(\bm{\tau}\cdot\partial_{\mu}\bm{\pi}+\frac{\partial_{\mu}{\eta}}{\sqrt{3}}\right)\right$. (17)
$+$ $\left.\left(\partial_{\mu}\bar{K}K^{*\mu}-\bar{K}^{*\mu}\partial_{\mu}K\right)\left(\bm{\tau}\cdot\bm{\pi}+\frac{\eta}{\sqrt{3}}\right)\right],$
$\mathcal{L}_{{V}KK}$ $=$ $\frac{ig}{4}\left[\bar{K}\partial_{\mu}K-\partial_{\mu}\bar{K}K\right]\left(\bm{\tau}\cdot\bm{\rho}^{\mu}+{\omega}^{\mu}\right),$ (18)
$\mathcal{L}_{{V}K^{*}K^{*}}$ $=$ $\frac{ig}{4}\left[\left(\bar{K}_{\mu}^{*}\partial^{\mu}K^{*\nu}-\partial^{\mu}\bar{K}^{*\nu}K_{\mu}^{*}\right)\left(\bm{\tau}\cdot\bm{\rho}_{\nu}+\omega_{\nu}\right)\right$. (19)
$\left.+\left(\partial^{\mu}\bar{K}^{*\nu}K_{\nu}^{*}-\bar{K}_{\nu}^{*}\partial^{\mu}K^{*\nu}\right)\left(\bm{\tau}\cdot\bm{\rho}_{\mu}+\omega_{\mu}\right)\right$.
$\left.+\left(\bar{K}_{\nu}^{*}K^{*}_{\mu}-\bar{K}_{\mu}^{*}K^{*}_{\nu}\right)\left(\bm{\tau}\cdot\partial^{\mu}\bm{\rho}^{\nu}+\partial^{\mu}\omega^{\nu}\right)\right],$
$\mathcal{L}_{PK^{*}K^{*}}$ $=$ $g_{VVP}\varepsilon_{\mu\nu\alpha\beta}\partial^{\mu}\bar{K}^{*\nu}\partial^{\alpha}K^{*\beta}\left(\bm{\tau}\cdot\bm{\pi}+\frac{\eta}{\sqrt{3}}\right),$ (20)
$\mathcal{L}_{VKK^{*}}$ $=$ $g_{VVP}\varepsilon_{\mu\nu\alpha\beta}\left(\partial^{\mu}\bar{K}^{*\nu}K+\bar{K}\partial^{\mu}{K}^{*\nu}\right)$ (21)
$\left(\bm{\tau}\cdot\partial^{\alpha}\bm{\rho}^{\beta}+\partial^{\alpha}{\omega}^{\beta}\right).$
Table 1: The effective potentials for $Y_{b}(\Lambda_{b},\Sigma_{b})\bar{K}^{(*)}$ systems.
Processes Effective potentials
$V_{\Lambda_{b}\bar{K}^{*}\to\Lambda_{b}\bar{K}^{*}}$ $l_{B}g_{\sigma}(\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{4}^{{\dagger}})\chi_{3}^{{\dagger}}\chi_{1}Y(\Lambda,m_{\sigma},r)-\frac{\beta_{B}g_{V}g}{4}(\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{4}^{{\dagger}})\chi_{3}^{{\dagger}}\chi_{1}Y(\Lambda,m_{\omega},r)$
$V_{\Lambda_{b}\bar{K}^{*}\to\Sigma_{b}\bar{K}^{*}}$ $\frac{1}{6}\frac{g_{4}g_{VVP}}{f_{\pi}}\mathcal{F}_{1}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})Y(\Lambda_{0},m_{\pi 0},r)-\frac{1}{6\sqrt{2}}\frac{\lambda_{I}g_{V}g}{m_{K^{*}}}\mathcal{F}_{2}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})Y(\Lambda_{0},m_{\rho 0},r)$
$V_{\Sigma_{b}\bar{K}^{*}\to\Sigma_{b}\bar{K}^{*}}$ $\frac{1}{2}l_{S}g_{\sigma}\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}Y(\Lambda,m_{\sigma},r)+\frac{g_{1}g_{VVP}}{6\sqrt{2}f_{\pi}}\mathcal{F}_{1}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})\mathcal{G}(I)Y(\Lambda,m_{\pi},r)-\frac{g_{1}g_{VVP}}{18\sqrt{2}f_{\pi}}\mathcal{F}_{1}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})Y(\Lambda,m_{\eta},r)$
$\frac{1}{8}\beta_{S}g_{V}g\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}\mathcal{G}(I)Y(\Lambda,m_{\rho},r)+\frac{\lambda_{S}g_{V}g}{8\sqrt{3}m_{\Sigma_{b}}}\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}\mathcal{G}(I)\nabla^{2}Y(\Lambda,m_{\rho},r)-\frac{\lambda_{S}g_{V}g}{24\sqrt{3}m_{K^{*}}}\mathcal{F}_{2}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})\mathcal{G}(I)Y(\Lambda,m_{\rho},r)$
$\frac{1}{8}\beta_{S}g_{V}g\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}Y(\Lambda,m_{\omega},r)-\frac{\lambda_{S}g_{V}g}{8\sqrt{3}m_{\Sigma_{b}}}\chi_{3}^{{\dagger}}\chi_{1}\bm{\epsilon}_{2}\cdot\bm{\epsilon}_{3}^{{\dagger}}\nabla^{2}Y(\Lambda,m_{\omega},r)+\frac{\lambda_{S}g_{V}g}{24\sqrt{3}m_{K^{*}}}\mathcal{F}_{2}(r,\bm{\sigma},i\bm{\epsilon}_{2}\times\bm{\epsilon}_{4}^{{\dagger}})Y(\Lambda,m_{\omega},r)$
$V_{\Sigma_{b}\bar{K}\to\Sigma_{b}\bar{K}}$ $\frac{1}{2}l_{S}g_{\sigma}\chi_{3}^{{\dagger}}\chi_{1}Y(\Lambda,m_{\sigma},r)+\frac{\mathcal{G}(I)}{8}\beta_{S}g_{V}g\chi_{3}^{{\dagger}}\chi_{1}Y(\Lambda,m_{\rho},r)-\frac{\mathcal{G}(I)}{24m_{\Sigma_{b}}}\lambda_{S}g_{V}g\chi_{3}^{{\dagger}}\chi_{1}\nabla^{2}Y(\Lambda,m_{\rho},r)$
$\frac{1}{8}\beta_{S}g_{V}g\chi_{3}^{{\dagger}}\chi_{1}Y(\Lambda,m_{\omega},r)+\frac{1}{24m_{\Sigma_{b}}}\lambda_{S}g_{V}g\chi_{3}^{{\dagger}}\chi_{1}\nabla^{2}Y(\Lambda,m_{\omega},r)$
$V_{\Lambda_{b}\bar{K}^{*}\to\Sigma_{b}\bar{K}}$ $\frac{1}{6}\frac{g_{4}g}{f_{\pi}\sqrt{m_{K}m_{K^{*}}}}\mathcal{F}_{1}(r,\bm{\sigma},\bm{\epsilon}_{2})U(\Lambda_{1},m_{\pi 1},r)-\frac{\lambda_{I}g_{V}g_{VVP}}{3\sqrt{2}}\sqrt{\frac{m_{K^{*}}}{m_{K}}}\mathcal{F}_{2}(r,\bm{\sigma},\bm{\epsilon}_{2})Y(\Lambda_{1},m_{\rho 1},r)$
$V_{\Sigma_{b}\bar{K}^{*}\to\Sigma_{b}\bar{K}}$ $\frac{g_{1}g\mathcal{F}_{1}(r,\bm{\sigma},\bm{\epsilon}_{2})}{24\sqrt{2}f_{\pi}\sqrt{m_{K}m_{K^{*}}}}\mathcal{G}(I)Y(\Lambda_{2},m_{\pi 2},r)+\frac{g_{1}g}{72\sqrt{2}f_{\pi}\sqrt{m_{K}m_{K^{*}}}}\mathcal{F}_{1}(r,\bm{\sigma},\bm{\epsilon}_{2})Y(\Lambda_{2},m_{\eta 2},r)$
$\frac{\lambda_{S}g_{V}g_{VVP}}{6\sqrt{3}}\sqrt{\frac{m_{K^{*}}}{m_{K}}}\mathcal{F}_{2}(r,\bm{\sigma},\bm{\epsilon}_{2})\mathcal{G}(I)Y(\Lambda_{2},m_{\rho 2},r)-\frac{\lambda_{S}g_{V}g_{VVP}}{6\sqrt{3}}\sqrt{\frac{m_{K^{*}}}{m_{K}}}\mathcal{F}_{2}(r,\bm{\sigma},\bm{\epsilon}_{2})Y(\Lambda_{2},m_{\omega 2},r)$
Before applying the potentials in coordinate space, we adopt the Breit approximation, a standard approach within the one-boson-exchange (OBE) framework. In this approximation, the relativistic scattering amplitude is reduced to an instantaneous effective potential in momentum space, which is valid in the low-energy regime with small momentum transfer. This treatment enables us to extract spin-dependent interaction terms, such as central, spin-spin, tensor, and spin-orbit components. For a more detailed discussion, we refer the reader to Refs. Breit:1929zz ; Breit:1930zza . Thus, the effective potential in momentum space reads
$\mathcal{V}^{h_{1}h_{2}\to h_{3}h_{4}}(\bm{q})$ $=$ $-\frac{\mathcal{M}(h_{1}h_{2}\to h_{3}h_{4})}{4\sqrt{m_{1}m_{2}m_{3}m_{4}}},$ (22)
in which $\mathcal{M}(h_{1}h_{2}\to h_{3}h_{4})$ denotes the scattering amplitude for the $h_{1}h_{2}\to h_{3}h_{4}$ process, and $m_{i}$ is the mass of the particle $h_{i}$. The Fourier transform with respect to $\bm{q}$ leads to the effective potential in position space,
$\mathcal{V}(\bm{r})=\int\frac{d^{3}\bm{q}}{(2\pi)^{3}}e^{i\bm{q}\cdot\bm{r}}\mathcal{V}(\bm{q})\mathcal{F}^{2}(q^{2},m_{i}^{2}),$ (23)
where $\mathcal{F}$ denotes the form factor with explicit form
$\mathcal{F}(q^{2},m_{i}^{2})=\frac{\Lambda^{2}-m_{i}^{2}}{\Lambda^{2}-q^{2}}.$ (24)
Here, the parameter $\Lambda$ is introduced as an UV cut-off originates from the fact that the hadrons have a non-zero size to account for the internal structure of the interacting hadrons.
The corresponding one-boson-exchange effective potentials are taken from Ref. Chen:2022svh and listed in Table 1, where $\mathcal{G}=-2$ for the $I=1/2$ system and $\mathcal{G}=1$ for the $I=3/2$ system. The explicit expressions for factors $\mathcal{F}_{1,2}$, $U$ and $Y$ in the effective potentials listed in Table 1 are given
$\mathcal{F}_{1}(r,\mathbf{a},\mathbf{b})$ $=$ $\chi_{3}^{\dagger}\bigg(\mathbf{a}\cdot\mathbf{b}\nabla^{2}+S(\hat{r},\mathbf{a},\mathbf{b})r\frac{\partial}{\partial r}\frac{1}{r}\frac{\partial}{\partial r}\bigg)\chi_{1}$
$\mathcal{F}_{2}(r,\mathbf{a},\mathbf{b})$ $=$ $\chi_{3}^{\dagger}\bigg(2\mathbf{a}\cdot\mathbf{b}\nabla^{2}-S(\hat{r},\mathbf{a},\mathbf{b})r\frac{\partial}{\partial r}\frac{1}{r}\frac{\partial}{\partial r}\bigg)\chi_{1}$
$U(\Lambda,m,r)$ $=$ $\frac{1}{4\pi r}\bigg(\text{cos}(mr)-e^{-\Lambda r}\bigg)-\frac{\Lambda^{2}+m^{2}}{8\pi\Lambda}e^{-\Lambda r}$
$Y(\Lambda,m,r)$ $=$ $\frac{1}{4\pi r}(e^{-mr}-e^{-\Lambda r})-\frac{\Lambda^{2}-m^{2}}{8\pi\Lambda}e^{-\Lambda r},$ (25)
where $S(\hat{r},\mathbf{a},\mathbf{b})\equiv 3(\hat{r}\cdot\mathbf{a})(\hat{r}\cdot\mathbf{b})-\mathbf{a}\cdot\mathbf{b}$. The values of relevant parameters are listed in Table 2 Liu:2011xc ; Chen:2017xat ; Kaymakcalan:1983qq . The subscript labels $m_{\pi 1}$, $m_{\pi 2}$, etc., represent the effective masses of the exchanged mesons, which arise from the energy transfer $q^{0}$ during the Fourier transformation. Specifically, for the $Y(\Lambda_{2},m_{\pi 2},r)$-type potential, we use
$m_{\pi 2}$ $=$ $\sqrt{m_{\pi}^{2}-(q^{0})^{2}},\quad\Lambda_{2}=\sqrt{\Lambda^{2}-(q^{0})^{2}},$ (26)
$q^{0}$ $=$ $\frac{m_{1}^{2}+m_{4}^{2}-m_{2}^{2}-m_{3}^{2}}{2(m_{3}+m_{4})}=\frac{m_{\bar{K}}^{2}-m_{\bar{K}^{*}}^{2}}{2(m_{\Sigma_{b}}+m_{\bar{K}})}.$ (27)
For the $U(\Lambda_{1},m_{\pi 1},r)$-type potential, the definitions are
$m_{\pi 1}$ $=$ $\sqrt{(q^{0})^{2}-m_{\pi}^{2}},\quad\Lambda_{1}=\sqrt{\Lambda^{2}-(q^{0})^{2}},$ (28)
$q^{0}$ $=$ $\frac{m_{1}^{2}+m_{4}^{2}-m_{2}^{2}-m_{3}^{2}}{2(m_{3}+m_{4})}=\frac{m_{\Lambda_{b}}^{2}+m_{\bar{K}}^{2}-m_{\bar{K}^{*}}^{2}-m_{\Sigma_{b}}^{2}}{2(m_{\Sigma_{b}}+m_{\bar{K}})}.$ (29)
The same conventions apply to other exchanged mesons such as $\rho_{1,2}$, $\eta_{2}$, and $\omega_{1,2}$.
Table 2: The relevant parameters adopted in this work.
Parameters Values
$l_{s}=2l_{B}$ 7.3
$g_{1}=(\sqrt{8}/3)g_{4}$ 1.0
$\beta_{s}g_{v}=-2\beta_{B}g_{v}$ 12.0
$\lambda_{s}g_{v}=-2\sqrt{2}\lambda_{I}g_{v}$ $19.2~\textrm{GeV}^{-1}$
$g_{\sigma}$ 3.65-3.65
$g$ 12.00
$g_{VVP}$ $3g^{2}/(32\sqrt{2}\pi^{2}f_{\pi})$
$f_{\pi}$ 0.132 GeV
II.2 Complex scaling method
At present, in order to obtain possible poles for these investigated systems, the complex scaling method (CSM) is applied. In the CSM, the relative distance $\bm{r}$ and the conjugate momentum $\bm{p}$ are replaced by
$\bm{r^{\prime}}\to\bm{r}e^{i\theta},\quad\bm{p^{\prime}}\to\bm{p}e^{-i\theta}$ (30)
where the scaling angle $\theta$ is chosen to be positive. Applying such a replacement to the Schrödinger equation, we get the complex-scaled Schrödinger equation for the coupled channels which reads
$\left[\frac{1}{2\mu_{j}}\left(-\frac{d^{2}}{dr^{2}}+\frac{l_{j}(l_{j}+1)}{r^{2}}\right)e^{-2i\theta}+W_{j}\right]\psi_{j}^{\theta}(r)$
$\qquad+\sum_{k}V_{jk}(re^{i\theta})\psi_{k}^{\theta}(r)=E\psi_{j}^{\theta}(r),$ (31)
where $\mu_{j}$, $W_{j}$, and $\psi^{\theta}_{j}(r)$ are the reduced mass, corresponding threshold, and the orbital wave function, respectively.
It is worth mentioning that the properties of the solutions of the complex scaling Schrödinger equation can be predicted by the so-called ABC theorem Aguilar:1971ve ; Balslev:1971vb , which means
  1. 1.
    The wave functions for resonant states should be square-integrable on the complex plane, which is the same as for a bound state.
  2. 2.
    On the complex plane, the eigenvalues of the bound states and resonances are independent of the scaling angle $\theta$.
  3. 3.
    The continuum states change along the $2\theta$ line.
Figure 2: Schematic eigenvalue distributions of $H_{\theta}$ in the coupled-channel two-body systems.
According to this theorem, one pick out can the bound state or resonance from the continuum states as shown in Figure 2. This technique has been widely and successfully applied in quantum scattering theory and few-body systems. In our work, it provides a robust tool for identifying resonances in the coupled channel systems. Moreover, in this work, the orbital wave functions are expanded in terms of a set of Gaussian basis functions. With the obtained wave functions, the root-mean-square (RMS) radii $r_{RMS}$ and component proportions $P$ can be calculated by T.yo ; Lin:2023ihj
$r_{RMS}^{2}$ $=\langle\psi^{\theta}|r^{2}|\psi^{\theta}\rangle=\sum_{i}\int r^{2}\psi^{\theta}_{i}(\bm{r})^{2}d^{3}\bm{r},$ (32a)
$P$ $=\langle\psi^{\theta}_{i}|\psi^{\theta}_{i}\rangle=\int\psi^{\theta}_{i}(\bm{r})^{2}d^{3}\bm{r},$ (32b)
where the $\psi^{\theta}_{i}$ are normalized as
$\sum_{i}\langle\psi^{\theta}_{i}\mid\psi^{\theta}_{i}\rangle=1.$ (33)
Interestingly, the root mean square (RMS) radius $r_{\rm RMS}$ for the resonances can be a complex number. In such cases, one can use the interpretation scheme proposed by T. Berggren, which generalizes the concept of expectation values from bound states to resonances Berggren:1970wto . According to this scheme, the real part of the complex $r_{\rm RMS}$ represents the usual physical expectation value, while the imaginary part indicates a measure of uncertainty in the observation. Numerical calculations of $r^{2}$ have supported this generalized interpretation Gyarmati:1972yac ; 1997matrix . Besides, the $P$ involves the square of the complex-scaled wave function $\psi^{\theta}_{i}(\bm{r})$, rather than its modulus squared. Consequently, $P$ is not strictly positive definite. Depending on the nodal structure and the phase of the wave function, $P$ can take on positive or negative real values, or even have a complex part. Moreover, it is worth mentioning that the scaling angle $\theta$ should be larger than $1/2\text{Arg}(\Gamma/2E)$ to ensure the normalizability of the wave functions of the resonant states Myo:2014ypa .
Figure 3: The $\Lambda$ dependence for the bottom-strange pentaquarks systems. The red solid dots stand for the bound states. The blue open circles with bars correspond to the resonances, with the lengths of bars being the total widths of the corresponding resonances.
III Results and discussions
Figure 4: The complex energy eigenvalues of $I(J^{P})=1/2(1/2^{-})$ system by varying the angle $\theta$ from $30^{\circ}\sim 40^{\circ}$.
Figure 5: (Color online) The $r$ dependence of the deduced effective potentials with cutoff of 1000 $\rm{MeV}$ for the $S$-wave $\Sigma_{b}\bar{K}^{*}$ system with $I(J^{P})=1/2(1/2^{-})$.
Performing the above procedure, we can systematically investigate the bottom-strange molecular pentaquarks by solving the coupled channel Schrödinger equation. In this work, the only free parameter is the UV cutoff $\Lambda$ in Eq. (24), which may vary for different coupled systems being investigated, and it lies within the range of $800\sim 5000$ MeV Liu:2009qhy ; Chen:2019asm . A reasonable cutoff for deuteron studies is estimated to be around 1000 $\rm MeV$, but it can also vary among different scenarios Tornqvist:1993ng ; Tornqvist:1993vu . A loosely bound state at a cutoff around $1000$ $\rm MeV$ could be a promising molecular state candidate; thus, we summarize our predictions for bottom-strange pentaquark molecular state systems with cutoff $\Lambda$ in a range of $800\sim 1100$ $\rm MeV$ in Table 5.
We firstly deal with the bottom baryon $Y_{b}$ and $\bar{K}^{*}$ meson systems to reveal possible bound and resonant states and give them reasonable interpretations. The same technique can be applied in the analysis of the bottom baryon $Y_{b}$ and $K$ meson systems. Our estimations for these investigated systems depending on the cutoff value $\Lambda$ are plotted in Figure 3 and listed in Table 3 and 4. In the present work, both $S$-$D$ wave mixing effects and coupled channel effects are taken into account. According to the isospin, spin, and parity, the bottom baryon and anti-strange meson systems can be classified as $1/2(1/2^{-})\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$, $3/2(1/2^{-})\Sigma_{b}\bar{K}/\Sigma_{b}\bar{K}^{*}$, $1/2(3/2^{-})\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$, and $3/2(3/2^{-})\Sigma_{b}\bar{K}^{*}$ channels, respectively. The corresponding classification also exists for the bottom baryon and strange meson systems.
Table 3: The numerical results for the obtained bound states.
$\Lambda(\rm{MeV})$ $r_{RMS}(\rm{fm})$ $E(\rm{MeV})$ $(\Sigma_{b}\bar{K}(^{2}S_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{2}S_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{4}D_{3/2}))$
$I(J^{P})=3/2(1/2^{-})$ 3600 3.77 6304.84 (99.08 0.47 0.45)
3850 1.44 6292.68 (95.75 2.28 1.97)
4100 1.03 6278.37 (92.78 3.97 3.25)
$\Lambda(\rm{MeV})$ $r_{RMS}(\rm{fm})$ $E(\rm{MeV})$ ($\Lambda_{b}\bar{K}^{*}(^{4}S_{3/2})$ $\Lambda_{b}\bar{K}^{*}(^{2}D_{3/2})$ $\Lambda_{b}\bar{K}^{*}(^{4}D_{3/2})$ $\Sigma_{b}\bar{K}^{*}(^{4}S_{3/2})$ $\Sigma_{b}\bar{K}^{*}(^{2}D_{3/2})$ $\Sigma_{b}\bar{K}^{*}(^{4}D_{3/2})$)
$I(J^{P})=1/2(3/2^{-})$ 1460 1.06 6511 (12.19 0.52 2.98 75.64 1.15 7.52)
1480 0.59 6493 (4.23 0.53 3.00 82.90 1.22 8.12)
1500 0.55 6474 (3.14 0.50 2.88 83.93 1.23 8.32)
$\Lambda(\rm{MeV})$ $r_{RMS}(\rm{fm})$ $E(\rm{MeV})$ ($\Sigma_{b}\bar{K}^{*}(^{4}S_{3/2})$ $\Sigma_{b}\bar{K}^{*}(^{2}D_{3/2})$ $\Sigma_{b}\bar{K}^{*}(^{4}D_{3/2})$)
$I(J^{P})=3/2(3/2^{-})$ 1300 3.90 6704 (99.26 0.16 0.58)
1400 2.36 6721 (98.67 0.29 1.04)
1500 1.61 6697 (98.17 0.40 1.43)
$\Lambda(\rm{MeV})$ $r_{RMS}(\rm{fm})$ $E(\rm{MeV})$ $(\Sigma_{b}K(^{2}S_{1/2})$ $\Sigma_{b}K^{*}(^{2}S_{1/2})$ $\Sigma_{b}K^{*}(^{4}D_{1/2}))$
$I(J^{P})=3/2(1/2^{-})$ 1260 1.51 6298.01 (76.17 23.72 0.10)
1270 0.90 6284.01 (65.82 34.04 0.15)
1280 0.67 6265.52 (58.84 41.00 0.16)
$\Lambda(\rm{MeV})$ $r_{RMS}(\rm{fm})$ $E(\rm{MeV})$ ($\Lambda_{b}K^{*}(^{4}S_{3/2})$ $\Lambda_{b}K^{*}(^{2}D_{3/2})$ $\Lambda_{b}K^{*}(^{4}D_{3/2})$ $\Sigma_{b}K^{*}(^{4}S_{3/2})$ $\Sigma_{b}K^{*}(^{2}D_{3/2})$ $\Sigma_{b}K^{*}(^{4}D_{3/2})$)
$I(J^{P})=1/2(3/2^{-})$ 1320 2.02 6510 (70.16 0.09 0.37 27.76 0.32 1.30)
1340 0.96 6502 (52.39 0.11 0.54 44.70 0.43 1.83)
1360 0.70 6490 (43.51 0.12 0.58 53.39 0.45 1.95)
Table 4: The numerical results for the bottom baryon and anti-meson interactions with $I(J^{P})=1/2(1/2^{-})$ .
$\Lambda(\rm{MeV})$ $r_{RMS}(\rm{fm})$ $E(\rm{MeV})$ ($(\Sigma_{b}\bar{K}^{*}(^{2}S_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{4}D_{1/2})$)
$I(J^{P})=1/2(1/2^{-})$ 840 4.37 6703.34 (99.65 0.35)
970 1.16 6687.43 (99.15 0.85)
1100 0.78 6651.48 (98.83 1.17)
$\Lambda(\rm{MeV})$ $r_{RMS}(\rm{fm})$ $E(\rm{MeV})$ ($(\Sigma_{b}\bar{K}(^{2}S_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{2}S_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{4}D_{1/2})$)
$I(J^{P})=1/2(1/2^{-})$ 900 1.890.20i1.89-0.20i 6699.720.96i6699.72-0.96i (0.150.13i-0.15-0.13i $99.39+0.13i$ 0.760.00i0.76-0.00i)
1000 $1.12+0.00i$ 6680.000.07i6680.00-0.07i ($0.02+0.04i$ 99.080.05i99.08-0.05i $0.90+0.01i$)
1100 0.820.03i0.82-0.03i 6643.7812.99i6643.78-12.99i (1.13+0.29i-1.13+0.29i 100.640.62i100.64-0.62i $0.49+0.33i$)
$\Lambda(\rm{MeV})$ $r_{RMS}(\rm{fm})$ $E(\rm{MeV})$ $($$\Lambda_{b}\bar{K}^{*}(^{2}S_{1/2})$ $\Lambda_{b}\bar{K}^{*}(^{4}D_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{2}S_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{4}D_{1/2})$)
$I(J^{P})=1/2(1/2^{-})$ 920 $2.46+5.02i$ 6697.4833.62i6697.48-33.62i (16.57+2.89i-16.57+2.89i $0.28+1.26i$ 116.264.14i116.26-4.14i 0.030.01i0.03-0.01i)
$10$ 1.010.22i1.01-0.22i 6651.1042.66i6651.10-42.66i (31.38+10.64i-31.38+10.64i $0.95+1.67i$ 130.0313.15i130.03-13.15i $0.40+0.84i$)
1100 0.880.27i0.88-0.27i 6605.0331.83i6605.03-31.83i (13.67+28.50i-13.67+28.50i $1.46+0.43i$ 110.8229.83i110.82-29.83i $1.39+0.90i$)
$\Lambda(\rm{MeV})$ $r_{RMS}(\rm{fm})$ $E(\rm{MeV})$ ($\Sigma_{b}\bar{K}(^{2}S_{1/2})$ $\Lambda_{b}\bar{K}^{*}(^{2}S_{1/2})$ $\Lambda_{b}\bar{K}^{*}(^{4}D_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{2}S_{1/2})$ $\Sigma_{b}\bar{K}^{*}(^{4}D_{1/2})$)
$I(J^{P})=1/2(1/2^{-})$ 960 1.220.55i1.22-0.55i 6692.8417.29i6692.84-17.29i ($1.24+0.25i$ 2.994.37i-2.99-4.37i 1.15+0.68i-1.15+0.68i $102.41+3.33i$ $0.49+0.11i$)
1030 0.980.09i0.98-0.09i 6659.0518.99i6659.05-18.99i ($1.36+0.35i$ 5.99+0.17i-5.99+0.17i 0.46+1.34i-0.46+1.34i 104.282.40i104.28-2.40i $0.81+0.54i$)
1100 0.840.10i0.84-0.10i 6615.8015.50i6615.80-15.50i ($0.89+0.51i$ 4.14+7.39i-4.14+7.39i $0.40+1.11i$ 101.579.75i101.57-9.75i $1.28+0.74i$)
III.1 Bottom baryon and anti-strange meson systems
III.1.1 $I(J^{P})=1/2(1/2^{-})$ with $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$
In this subsection, we firstly discuss the coupled $I(J^{PC})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ systems. As shown in Fig. 3(a), one could find that a bound state and a resonance emerge within the range $\Lambda=880\sim 1100$ $\rm{MeV}$. When the cutoff $\Lambda$ is set to 880 $\rm{MeV}$, the bound state is located below the $\Sigma_{b}\bar{K}$ threshold with a binding energy of about 44 $\rm{MeV}$, the $r_{RMS}$ is 3 $\rm{fm}$, and it is dominated by the $\Sigma_{b}\bar{K}(^{2}S_{1/2})$ channel. Sliding the cutoff to $1080$ $\rm{MeV}$, the mass becomes around 6222 $\rm{MeV}$ and the $r_{RMS}$ is 0.8 $\rm{fm}$, which is consistent with the sizes of exotic hadronic molecular states. Thus, this bound state is a good candidate for the particle $\Xi_{b}(6227)$. These results favor the conclusion in Refs. Huang:2018bed ; Zhu:2020lza ; Mutuk:2024elj .
Meanwhile, at a cutoff of $\Lambda=960~\rm{MeV}$, we obtain a quasi-bound state with complex energy $E=6693-17i~\rm{MeV}$ and root-mean-square radius $r_{\rm RMS}=1.22-0.55i~\rm{fm}$, which is predominantly composed of the $\Sigma_{b}\bar{K}^{*}(^{2}S_{1/2})$ channel. The binding of this state mainly arises from the strong interaction in the (closed) $\Sigma_{b}\bar{K}^{*}$ channel, while the nonzero width is induced via its coupling to lower open channels such as $\Lambda_{b}\bar{K}^{*}$. This structure is consistent with the nature of a Feshbach-type resonance and can be regarded as a good hadronic molecular candidate.
Besides, We illustrate the complex energy eigenvalues of the $I(J^{P})=1/2(1/2^{-})$ system while varying the angle $\theta$ from $30^{\circ}\sim 40^{\circ}$ in Figure 4. Specifically, the complex energy plane reveals the nature of poles: bound states lie on the real axis below threshold, while resonance appears as pole in the complex plane. The behavior of continuum states under varying scaling angle $\theta$ illustrates the separation between physical and unphysical states. Additionally, the clustering of poles near certain thresholds, such as $\Sigma_{b}\bar{K}$, $\Lambda_{b}\bar{K}^{*}$, and $\Sigma_{b}\bar{K}^{*}$ highlights the role of channel coupling and threshold effects. These patterns help identify the dominant configurations and offer insight into the underlying interaction dynamics.
Furthermore, our results highlight that the pion-exchange potential plays a crucial role in forming this resonance, whereas the contribution from the $\eta$-exchange interaction is negligible. As shown in Fig. 5, the $\eta$ exchange provides a much weaker attractive potential compared to the $\pi$ exchange. The dominant role of the pion exchange arises from its lighter mass and longer range, which enhance the attractive interaction between hadrons. In contrast, the $\eta$ meson is heavier, and its contribution is significantly suppressed. This physical picture is now clearly stated in the revised text. Moreover, according to our calculations, the resonance pole still appears even when the $\eta$-exchange potential is omitted, provided that the same cutoff is used.
III.1.2 $I(J^{P})=3/2(1/2^{-})$ with $\Sigma_{b}\bar{K}/\Sigma_{b}\bar{K}^{*}$
[…]
III.1.3 $I(J^{P})=1/2(3/2^{-})$ with $\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$
Besides, we investigate the $I(J^{P})=1/2(3/2^{-})$ $\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ channel for the $Y_{b}$ and $\bar{K}^{(*)}$ system, and the corresponding results are listed in Table 3. According to our estimations, at a cutoff of 1460 MeV, a bound state appears below the $\Lambda_{b}\bar{K}^{*}$ threshold. This state is mainly composed of the $\Sigma_{b}\bar{K}^{*}(^{4}S_{3/2})$ (75.6%), $\Sigma_{b}\bar{K}^{*}(^{4}D_{3/2})$ (7.5%), and $\Lambda_{b}\bar{K}^{*}(^{4}S_{3/2})$ (12%) components, and shows a strong dependence on the cutoff value. However, since this cutoff is inconsistent with the empirical value used for describing deuteron systems, and the corresponding $r_{\rm RMS}$ does not fall within the typical range for a molecular state, we conclude that the $I(J^{P})=1/2(3/2^{-})$ $\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ system is unlikely to form a hadronic molecular state.
III.1.4 $I(J^{P})=3/2(3/2^{-})$ with $\Sigma_{b}\bar{K}^{*}$
In the $I(J^{P})=3/2(3/2^{-})$ $\Sigma_{b}\bar{K}^{*}$ system, a weakly bound state with an energy of $6703$ $\rm{MeV}$ appears at a cutoff of $1300$ $\rm{MeV}$ and is dominated by the $S$-wave channel about 99%. If only the one-pion-exchange potential is considered, a bound state is obtained with a cutoff of $2400$ $\rm{MeV}$, which means that the potentials of $\rho$ and $\omega$ exchanges are helpful to form the bound state.
III.2 Bottom baryon and strange meson systems
For $Y_{b}K^{(*)}$ systems, the effective potentials from the $\omega$ and $\pi$ exchanges are in complete contrast with those of the $Y_{b}\bar{K}^{(*)}$ systems. Unlike the bottom baryon and anti-strange meson systems, for the bottom baryon and strange meson systems, one can only obtain bound state solutions. The corresponding numerical results are collected in Table 3 and Figure 3(b).
III.2.1 $I(J^{P})=1/2(1/2^{-})$ with $\Sigma_{b}K/\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$
When the cutoff parameter varies within $920\sim 1100~\rm{MeV}$, the mass decreases from 6304 to 6269 $\rm{MeV}$, while the root-mean-square radius $r_{RMS}$ shrinks from 3 to 1 $\rm{fm}$, consistent with expectations for a molecular candidate. Considering only the single $\Sigma_{b}K$ channel, a bound state solution also emerges at a larger cutoff of $\Lambda=1900~\rm{MeV}$, further indicating that coupled-channel effects play a significant role in forming the molecule.
III.2.2 $I(J^{P})=3/2(1/2^{-})$ with $\Sigma_{b}K/\Sigma_{b}K^{*}$
For the $I(J^{P})=3/2(1/2^{-})$ $\Sigma_{b}K/\Sigma_{b}K^{*}$ system, a bound state solution is found when the cutoff lies in the $1260~\rm MeV$, which is mainly composed of $S$- wave $\Sigma_{b}K$ and $\Sigma_{b}K^{*}$ about $76\%$ and 24%, respectively. When we vary the cutoff to $1270~\rm MeV$, the predicted mass varies from 6298 to 6284 $\rm{MeV}$, with the corresponding $r_{RMS}$ decreasing from $1.5$ to $0.9~\rm{fm}$. This sensitivity to the cutoff suggests the possibility of a molecular state.
III.2.3 $I(J^{P})=1/2(3/2^{-})$ with $\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$
As listed in Table 3, for the $\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ system with $I(J^{P})=1/2(3/2^{-})$, a bound state below the $\Lambda_{b}K^{*}$ threshold appears at a cutoff of 1260 $\rm{MeV}$. However, if only the $\Lambda_{b}K^{*}$ channel is considered, no bound state pole is found for cutoff values ranging from 800 to 5000 $\rm MeV$. In contrast, a loosely bound state dominated by the $\Sigma_{b}K^{*}(^{4}S_{3/2})$ component (about 98%) appears when considering only the $\Sigma_{b}K^{*}$ channel, with a cutoff of 960 $\rm MeV$. This state has a mass of 6704 $\rm MeV$ and a root-mean-square radius $r_{\rm RMS}=4~\rm fm$.
III.2.4 $I(J^{P})=3/2(3/2^{-})$ with $\Sigma_{b}K^{*}$
Finally, for the $I(J^{P})=3/2(3/2^{-})$ $\Sigma_{b}K^{*}$ system, no bound state solution is found within the cutoff range $\Lambda=800\sim 5000~\rm{MeV}$.
[…]
$\frac{1}{2}(\frac{1}{2}^{-})$$\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ $6301\sim 6222$ - $2.49\sim 0.73$ $\checkmark$ $\Lambda_{b}\bar{K}/\Xi_{b}^{(\prime)}\pi$
$6693\sim 6516$ $34.58\sim 31.00$ $1.22-0.55i\sim 0.84-0.10i$ $\checkmark$ $\Lambda_{b}\bar{K}^{(*)}/\Sigma_{b}\bar{K}/\Lambda\bar{B}^{(*)}/\Sigma\bar{B}/\Xi_{b}^{(\prime)}\pi/\Xi_{b}^{(\prime)}\eta/\Xi_{b}\rho/\Xi_{b}\omega$
$\frac{3}{2}(\frac{1}{2}^{-})$$\Sigma_{b}\bar{K}/\Sigma_{b}\bar{K}^{*}$ - - - $\times$ -
$\frac{1}{2}(\frac{3}{2}^{-})$$\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ - - - $\times$ -
$\frac{3}{2}(\frac{3}{2}^{-})\Sigma_{b}\bar{K}^{*}$ $6703\sim 6702$ - $3.37\sim 2.62$ $\checkmark$ $\Lambda_{b}\bar{K}^{*}/\Sigma_{b}^{*}\bar{K}/\Lambda\bar{B}^{*}/\Sigma\bar{B}^{*}/\Xi_{b}^{*}\pi/\Xi_{b}^{*}\eta/\Xi_{b}\rho/\Xi_{b}\omega$
$\frac{1}{2}(\frac{1}{2}^{-})$$\Sigma_{b}K/\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ $6303\sim 6269$ - $3.13\sim 1.08$ $\checkmark$ $N\bar{B}_{s}/\Lambda_{b}K$
$\frac{3}{2}(\frac{1}{2}^{-})$$\Sigma_{b}K/\Sigma_{b}K^{*}$ - - - $\times$ -
$\frac{1}{2}(\frac{3}{2}^{-})$$\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ - - - $\times$ -
$\frac{3}{2}(\frac{3}{2}^{-})$$\Sigma_{b}K^{*}$ $6704\sim 6692$ - $4.13\sim 1.34$ $\checkmark$ $N\bar{B}_{s}^{*}/\Lambda_{b}^{*}K/\Sigma_{b}^{*}K$
III.3 Further discussions
For $Y_{b}\bar{K}^{(*)}$ systems, we can obtain bound states and resonances, but only bound states are revealed for $Y_{b}K^{(*)}$ systems. The reason is that the flavor factors in the potentials for these systems differ significantly, which determine the relative sign and strength and are crucial for the formation of molecular states.
In our study, we aim to explore the crucial role of coupled-channel effects in the production of resonant states. As shown in Table 4, when considering only the single-channel $\Sigma_{b}\bar{K}^{*}$ $S$-$D$ wave mixing, we obtain a bound state at a cutoff of 840 $\rm{MeV}$. However, when including the coupled-channel effect for the $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Sigma_{b}\bar{K}^{*}$ system, we find a relatively narrow resonant state at a cutoff of 900 $\rm{MeV}$, which is dominated by the $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}^{*}(^{2}S_{1/2})$ component. Additionally, for the coupled-channel system $I(J^{P})=1/2(1/2^{-})$ $\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$, we obtain a relatively broad resonance at a cutoff of 920 $\rm{MeV}$. Furthermore, when considering the fully coupled $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ system, a resonance emerges near the $\Sigma_{b}\bar{K}^{*}$ threshold, with a complex energy $E{-}i\Gamma/2=6692.84{-}17.29i$ at a cutoff of 960 $\rm{MeV}$. On the other hand, when we discard the $D$-wave contributions, we still obtain a resonant state at a cutoff of approximately 1000 $\rm{MeV}$ for the $S$-wave coupled systems $\Sigma_{b}\bar{K}/\Sigma_{b}\bar{K}^{*}$, $\Lambda_{b}\bar{K}{/}\Sigma_{b}\bar{K}^{*}$, and $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ with $I(J^{P})=1/2(1/2^{-})$, respectively. In conclusion, our calculations demonstrate that the coupled-channel effect plays a significant role in the formation of resonant states, with the main contributions coming from the $S$-wave.
According to the masses and quantum numbers, we present some possible decay channels for these predicted states in Table 5. For instance, the $I(J^{PC})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ bound state can be found in the $\Lambda_{b}\bar{K}$ and $\Xi_{b}^{(\prime)}\pi$ channels. In the literature Chen:2018orb ; He:2021xrh ; Wang:2018fjm ; Cui:2019dzj , both $J^{P}=\frac{1}{2}^{-}$ molecular and $J^{P}=\frac{3}{2}^{-}/\frac{5}{2}^{-}$ conventional interpretations exist for the particle $\Xi_{b}(6227)$, and the spin is certainly crucial for distinguishing these two explanations. Another way to solve this puzzle is to hunt for the flavor exotic state with mass $6303\sim 6269~\rm{MeV}$ in the $I(J^{PC})=1/2(1/2^{-})$ $\Sigma_{b}K/\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ system, which is the mirror state of $\Xi_{b}(6227)$ in the molecular picture but does not appear in the three-quark picture. We strongly hope that future experiments can verify our proposals.
IV Summary
In this work, we systematically investigate the coupled $Y_{b}\bar{K}^{(*)}(Y_{b}K^{(*)})$ system to search for possible bound states and resonances by adopting the one-boson-exchange model within the complex scaling method. For the coupled $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ systems, according to our estimations, a bound state solution is obtained, which may correspond to the observed particle $\Xi_{b}(6227)$. Meanwhile, we find a $I(J^{PC})=1/2(1/2^{-})$ resonance near the $\Sigma_{b}\bar{K}^{*}$ threshold and a bound state in the $I(J^{P})=1/2(3/2^{-})$ $\Sigma_{b}\bar{K}^{*}$ system.
Then, when we extend our study to the $Y_{b}K^{(*)}$ systems, two loosely bound states are obtained. It is worth pointing out that the predicted bound state with mass $6303\sim 6269~\rm{MeV}$ in the $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}K/\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ system is flavor exotic and does not appear in the spectroscopy of conventional baryons, which provides a practical way to resolve the puzzle of the particle $\Xi_{b}(6227)$. We hope our predictions can offer valuable information to future experimental observations.
ACKNOWLEDGMENTS
We would like to thank Rui Chen, Xian-Hui Zhong, Li-Cheng Sheng, and Jin-Yu Huo for useful discussions. The work of X.-N. X. and Q. F. Song is supported by the National Natural Science Foundation of China under Grants No. 12275364. Q.-F. Lü is supported by the Natural Science Foundation of Hunan Province under Grant No. 2023JJ40421, the Scientific Research Foundation of Hunan Provincial Education Department under Grant No. 24B0063, and the Youth Talent Support Program of Hunan Normal University under Grant No. 2024QNTJ14.